# On relaxing the $N$-Reachability Implicit Requirement in NMPC Design

### v1

@@ 5 Illustrative example
where $\mu=0.4$ is a fixed parameter. $y_{1}$, $y_{2}$, $\theta$ are the planar coordinates and the roll angle to be controlled around some desired set-point. Denoting by $x=(y_{1},y_{2},\theta,\dot{y}_{1},\dot{y}_{2},\dot{\theta})\in\mathbb{R}^{6}$ the state vector. The setpoint is denoted by $x^{\text{ref}}=(y_{1}^{\text{ref}},y_{2}^{\text{ref}},0,0,0,0)$. The implementation uses a sampling period of $\tau=0.1$ and a prediction horizon $N=15$. Consider the following stage cost expressing this objective:
$\ell(x):=\|x-x^{\text{ref}}\|_{Q}^{2}+\|u-u^{d}\|_{R}^{2}$ (41)
(b) Without penalty on the final derivativeRefer to caption
Figure 2: Closed-loop evolution under the different settings and targeted states.
where $Q=\texttt{diag}(10^{2},10,10,1,1,1)$ and $R=\mathbb{I}_{2\times 2}$. We also consider a terminal penalty of the form $\|x-x^{\text{ref}}\|_{Q_{f}}^{2}$ with $Q_{f}=100\times Q$ in order to enhance the stability. This defines the standard stage cost and its associated standard penalty to which we might add the term $\Phi_{\gamma}$ introduced in the present contribution. This leads to the following cost function:
$\Bigl{[}\gamma^{2}\|f_{c}(x_{N},u_{N})\|+\gamma\ell(x_{N})\Bigr{]}+\underbrace{\texttt{standard cost}}_{\Psi_{f}(x_{N})+\sum_{i=0}^{N}\ell(x_{i},u_{i})}$ (42)
@@
Figure 2 shows that when the standard $N$-reachability condition is satisfied $(d=0.2)$, all settings leads to successful behavior. When this condition is strongly violated ($d\geq 1$), only the proposed formulation with sufficiently high $\gamma$ are successful (see the last two plots on Figure 2.a). In particular, it is useless to increase $\gamma$ while no penalty on the final derivative is applied as suggested by the last two plots of Figure 2.b. When standard reachability condition is almost satisfied ($d=0.5$), the setting without the derivative ($\gamma^{2}$)-related penalty might be successful provided that $\gamma$ is not too high. Finally Figure 3 shows that the use of high values of $\gamma$ (except for $\gamma=5000$) does not have detrimental effect on the computation time if not the contrary.

### v2

@@ 5 Illustrative example
where $\mu=0.4$ is a fixed parameter. $y_{1}$, $y_{2}$, $\theta$ are the planar coordinates and the roll angle to be controlled around some desired set-point. Denoting by $x=(y_{1},y_{2},\theta,\dot{y}_{1},\dot{y}_{2},\dot{\theta})\in\mathbb{R}^{6}$ the state vector. The set-point is denoted by $x^{\text{ref}}=(y_{1}^{\text{ref}},y_{2}^{\text{ref}},0,0,0,0)$. The implementation uses a sampling period of $\tau=0.1$ and a prediction horizon $N=15$. Consider the following stage cost expressing this objective:
$\ell(z):=\|x-x^{\text{ref}}\|_{Q}^{2}+\|u-u^{d}\|_{R}^{2}$ (41)
where $Q=\texttt{diag}(10^{2},10,10,1,1,1)$ and $R=\mathbb{I}_{2\times 2}$. We also consider a terminal penalty of the form $\|x-x^{\text{ref}}\|_{Q_{f}}^{2}$ with $Q_{f}=100\times Q$ in order to enhance the stability of the conventional formulation only. This defines the standard stage cost and its associated standard penalty to which we might add the term $\Phi_{\gamma}$ introduced in the present contribution. This leads to the following cost function ($\delta_{0,\gamma}$ is the Kronecker notation):
$\Bigl{[}\gamma^{2}\|f_{c}(x_{N},u_{N})\|+\gamma\ell(x_{N})\Bigr{]}+\underbrace{\texttt{standard cost}}_{\delta_{0,\gamma}\cdot\Psi_{f}(x_{N})+\sum_{i=0}^{N}\ell(x_{i},u_{i})}$ (42)
@@
(b) Without penalty on the final derivativeRefer to caption
Figure 2: Closed-loop evolution of the open-loop optimal cost $J^{\star}(z_{k}^{\text{cl}})$ under the different settings and targeted states.
@@
Figure 2 shows that when the standard $N$-reachability condition is satisfied $(d=0.2)$, all settings lead to successful behaviour. When this condition is strongly violated ($d\geq 1$), only the proposed formulation with sufficiently high $\gamma$ are successful (see the last plots on Figure 2.a). In particular, it is useless to increase $\gamma$ while no penalty on the final derivative is applied as suggested by the last two plots of Figure 2.b. When standard reachability condition is almost satisfied ($d=0.5$), the setting without the derivative ($\gamma^{2}$)-related penalty might be successful provided that $\gamma$ is not too high. Finally Figure 3 shows that the use of high values of $\gamma$ (except for $\gamma=5000$) does not have detrimental effect on the computation time if not the contrary.